X is a normally distributed random variable with a mean of 60 and standard






deviation of 8. Find the probabilities indicated by using the table.






(a) P(X < 52)






(b) P(48 < X < 64)






(c) P(X > 57)

1
Expert's answer
2022-03-29T10:59:28-0400

a. "P(X<52)=P(X<\frac{x-\mu}{\sigma})=P(X<\frac{52-60}{8})=P(X<-1)"

P(X<52)=0.1581

b.

"P(48<X<64)=P(\frac{48-64}{8}<X<\frac{64-60}{8})=P(-1.5<X<0.5)=0.1915+0.4332=0.6247"

c. "P(X>57)=P(X>\frac{57-60}{8})=P(X>-0.375)=1-P(X<-0.375)=1-0.35569=0.64431"